OpenStudy (anonymous):

how to integrate sin^2x-cos^2x?

OpenStudy (anonymous):

isn't this something like \(\cos(2x)\) ?

OpenStudy (anonymous):

or maybe its negative, one or the other

OpenStudy (nincompoop):

integrate (sin^2(x)-cos^2(x)) dx ?

OpenStudy (anonymous):

In solving an equation for x, you distribute a –4 on one side of the equation. What needs to be done to the other side of the equation? A. multiply by 4 B. add –4 C. subtract 4 D. nothing

OpenStudy (anonymous):

@nincompoop yes

OpenStudy (anonymous):

\[-\int\cos(2x)dx\]

OpenStudy (anonymous):

since \(\cos(2x)=\cos^2(x)-\sin^2(x)\) then a simple \(u\) sub will do it

OpenStudy (anonymous):

Yeah, use satellite's suggestion.